# Inference for Deep Neural Network Estimators in Generalized Nonparametric Models

### v3

@@ C.1 Preliminaries
$\Delta_{2}=\underbrace{\frac{1}{B^{2}}\sum_{i=1}^{n}\sum_{j=1}^{B}\big(\varepsilon_{b_{j}i}^{2}\big)-B^{-1}\mathrm{Var}_{i}}_{\Delta_{21}}+\underbrace{\frac{1}{B^{2}}\sum_{i=1}^{n}\sum_{j\neq j^{\prime}}\varepsilon_{b_{j}i}\varepsilon_{b_{j^{\prime}}i}}_{\Delta_{22}}.$
We have $\mathbb{E}_{b}\Delta_{21}=\mathbb{E}_{b}\Delta_{22}=0$. Simple calculation based on the randomness of subsampling gives $\mathrm{Var}_{b}(\Delta_{21})=O(n^{2}/B^{3})$, we have $|\Delta_{21}|=O_{p}(n/B^{3/2})$. As for $\Delta_{22}$, we consider $T_{b_{j},b_{j^{\prime}}}=\sum_{i=1}^{n}\varepsilon_{b_{j}i}\varepsilon_{b_{j^{\prime}}i}$, and we have $\mathbb{E}_{b}T_{b_{j},b_{j^{\prime}}}^{2}=O(r^{2}/n)$, and $\mathbb{E}_{b}T_{b_{j},b_{j^{\prime}}}\cdot T_{b^{\prime}_{j},b^{\prime}_{j^{\prime}}}=0$ as long as $b_{j},b_{j^{\prime}},b^{\prime}_{j},b^{\prime}_{j^{\prime}}$ has three or more distinct values by the conditional independence. We have
$\mathrm{Var}_{b}(\Delta_{22})\leq\frac{1}{B^{4}}\cdot\mathbb{E}_{b}\sum_{b_{j}\neq b_{j^{\prime}},b^{\prime}_{j}\neq b^{\prime}_{j^{\prime}}}T_{b_{j},b_{j^{\prime}}}\cdot T_{b^{\prime}_{j},b^{\prime}_{j^{\prime}}}=O\bigg(\frac{r^{2}}{nB^{2}}\bigg)$
@@
For the term $\Delta_{3}$, by the definition of $\widehat{\mathrm{Var}}_{i}=\frac{1}{B-1}\sum_{j=1}^{B}\bigl(Z_{b_{j}i}-\widehat{V}_{i}\bigr)^{2}$, it holds that
$\Delta_{3}$ $=\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}\bigg((Z_{b_{j}i}-\widehat{V}_{i})^{2}-\frac{B-1}{B}\mathrm{Var}_{i}\bigg)$
$=\bigg(\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}\bigg((Z_{b_{j}i}-M_{i})^{2}-\mathrm{Var}_{i}\bigg)\bigg)+O\bigg(\frac{n}{B^{2}}\bigg).$
Here, the first equality is by direct calculation and the second equality is by $\mathrm{Var}_{i}$ is bounded. Let , easy to bound that $\mathrm{Var}_{b}(T_{b_{j}})=O(n^{2})$, we have that
$\mathbb{E}_{b}\bigg\{\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}\bigg((Z_{b_{j}i}-M_{i})^{2}-\mathrm{Var}_{i}\bigg)\bigg\}=0;$
$\mathrm{Var}_{b}\bigg\{\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}\bigg((Z_{b_{j}i}-M_{i})^{2}-\mathrm{Var}_{i}\bigg)\bigg\}=O(n^{2}/B^{3}),$
@@
$\mathbb{E}(M_{i}-A_{i})^{2}$ $=\frac{r^{2}}{n^{2}}\sum_{k=1}^{r-1}{\binom{n-1}{k}}\bigg(\frac{{\binom{r-1}{k}}}{{\binom{n-1}{k}}}-\frac{{\binom{r}{k+1}}}{{\binom{n}{k+1}}}\bigg)^{2}V_{k+1}+\frac{r^{2}}{n^{2}}\sum_{k=2}^{r}{\binom{n-1}{k}}\bigg(\frac{{\binom{r-1}{k}}}{{\binom{n-1}{k}}}-\frac{{\binom{r}{k}}}{{\binom{n}{k}}}\bigg)^{2}V_{k}$
$\precsim\frac{r^{2}(n-1)}{n^{2}}\bigg(\frac{{r-1}}{{n-1}}-\bigg(\frac{{\binom{r}{2}}}{{\binom{n}{2}}}\bigg)\bigg)^{2}\binom{r}{2}^{-1}\mathrm{Var}[\widehat{f}^{b}(\mathbf{x}_{*})]$
$\precsim\frac{2r^{2}\mathrm{Var}[\widehat{f}^{b}(\mathbf{x}_{*})]}{n^{3}}.$ (C.10)
Here, the first inequality is by the fact that summation in the first line is maximized in the second-order terms and $\sum_{k=1}^{r}{\binom{r}{k}}V_{k}=\mathrm{Var}(\widehat{f}^{b}(\mathbf{x}_{*}))\leq C<+\infty$, and the second inequality is by simple calculation. We hence have
@@ Appendix D Proofs in Section C
$\mathbb{E}(Y|\eta)=\psi^{\prime}(\eta),\quad\mathrm{Var}(Y|\eta)=\psi^{\prime\prime}(\eta)\geq 0.$
We conclude that $\psi(\cdot)$ is a convex function. By Assumption 3.1, the non-degeneracy and the existence of distinct values in the support of $p(y|\eta)=h(y)\exp(\eta y-\psi(\eta))$ ensure that $\mathrm{Var}(Y|\eta)>0$ for all $\eta\in[-C,C]$. This guarantees that the second derivative of the log-partition function, $\psi^{\prime\prime}(\eta)=\mathrm{Var}(Y|\eta)$, is strictly positive. Furthermore, since $\mathbb{E}(Y|\eta)$ and $\mathbb{E}(Y^{2}|\eta)$ are defined as smooth integrals with respect to $p(y|\eta)$, they are continuous in $\eta$. As a result, $\psi^{\prime\prime}(\eta)$ is also continuous. By the compactness of $\eta\in[-C,C]$, there exists a constant $c>0$ such that:
@@
Lemma D.1 establishes the differences between $R_{n}(\widehat{f}_{n},f_{0})$ and $\widehat{R}_{n}(\widehat{f}_{n},f_{0})$. We now propose the following lemma, which addresses the challenges arising from the dependency of $\varepsilon_{i}$ on the covariates $\mathbf{x}_{i}$. This approach is different from previous work, such as Schmidt-Hieber (2020) and Fan & Gu (2024), which assumed independence between $\varepsilon_{i}$ and $\mathbf{x}_{i}$.
@@
$\mathrm{Var}\big(g_{j}(\bm{X}^{\prime}_{i},\mathbf{x}_{i})\big)$ $=2\mathrm{Var}(\ell(\bm{X};f_{j},f_{0}))$
@@
$\mathbb{P}(|\sum_{i=1}^{n}\xi_{i}|\geq t_{0})\leq 2\exp\bigg(-\frac{t_{0}^{2}}{2Mt_{0}/3+2\sum_{i=1}^{n}\mathrm{Var}(\xi_{i})}\bigg).$
@@
$\mathrm{Var}(\xi_{j}|(\mathbf{x}_{i})_{i})$ $=\frac{\sum_{i=1}^{n}\mathrm{Var}\big(\varepsilon_{i}(f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i}))|(\mathbf{x}_{i})_{i}\big)}{r_{j}^{2}}$
@@
where the first inequality is by $\mathrm{Var}(U|X)\leq\mathbb{E}(U^{2}|X)$ for all random variables $X$ and $U$, and the second inequality is by the definition of $r_{j}=\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\vee\sqrt{\sum_{i=1}^{n}(f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i}))^{2}/n}$, and $\mathbb{E}\big(\varepsilon_{i}^{2}|(\mathbf{x}_{i})_{i}\big)\leq 2\kappa^{2}$. Also, letting $a_{ij}=[f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i})]/r_{j}$, we have

### v4

@@ C.1 Preliminaries
$\Delta_{2}=\underbrace{\frac{1}{B^{2}}\sum_{i=1}^{n}\sum_{j=1}^{B}\big(\varepsilon_{b_{j}i}^{2}\big)-B^{-1}\operatorname{{\rm Var}}_{i}}_{\Delta_{21}}+\underbrace{\frac{1}{B^{2}}\sum_{i=1}^{n}\sum_{j\neq j^{\prime}}\varepsilon_{b_{j}i}\varepsilon_{b_{j^{\prime}}i}}_{\Delta_{22}}.$
We have $\mathbb{E}_{b}\Delta_{21}=\mathbb{E}_{b}\Delta_{22}=0$. Simple calculation based on the randomness of subsampling gives $\operatorname{{\rm Var}}_{b}(\Delta_{21})=O(n^{2}/B^{3})$, we have $|\Delta_{21}|=O_{p}(n/B^{3/2})$. As for $\Delta_{22}$, we consider $T_{b_{j},b_{j^{\prime}}}=\sum_{i=1}^{n}\varepsilon_{b_{j}i}\varepsilon_{b_{j^{\prime}}i}$, and we have $\mathbb{E}_{b}T_{b_{j},b_{j^{\prime}}}^{2}=O(r^{2}/n)$, and $\mathbb{E}_{b}T_{b_{j},b_{j^{\prime}}}\cdot T_{b^{\prime}_{j},b^{\prime}_{j^{\prime}}}=0$ as long as $b_{j},b_{j^{\prime}},b^{\prime}_{j},b^{\prime}_{j^{\prime}}$ has three or more distinct values by the conditional independence. We have
$\operatorname{{\rm Var}}_{b}(\Delta_{22})\leq\frac{1}{B^{4}}\cdot\mathbb{E}_{b}\sum_{b_{j}\neq b_{j^{\prime}},b^{\prime}_{j}\neq b^{\prime}_{j^{\prime}}}T_{b_{j},b_{j^{\prime}}}\cdot T_{b^{\prime}_{j},b^{\prime}_{j^{\prime}}}=O\bigg(\frac{r^{2}}{nB^{2}}\bigg)$
@@
For the term $\Delta_{3}$, by the definition of $\widehat{\operatorname{{\rm Var}}}_{i}=\frac{1}{B-1}\sum_{j=1}^{B}\bigl(Z_{b_{j}i}-\widehat{V}_{i}\bigr)^{2}$, it holds that
$\Delta_{3}$ $=\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}\bigg((Z_{b_{j}i}-\widehat{V}_{i})^{2}-\frac{B-1}{B}\operatorname{{\rm Var}}_{i}\bigg)$
$=\bigg(\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}\bigg((Z_{b_{j}i}-M_{i})^{2}-\operatorname{{\rm Var}}_{i}\bigg)\bigg)+O\bigg(\frac{n}{B^{2}}\bigg).$
Here, the first equality is by direct calculation and the second equality is by $\operatorname{{\rm Var}}_{i}$ is bounded. Let , easy to bound that $\operatorname{{\rm Var}}_{b}(T_{b_{j}})=O(n^{2})$, we have that
$\mathbb{E}_{b}\bigg\{\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}\bigg((Z_{b_{j}i}-M_{i})^{2}-\operatorname{{\rm Var}}_{i}\bigg)\bigg\}=0;$
$\operatorname{{\rm Var}}_{b}\bigg\{\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}\bigg((Z_{b_{j}i}-M_{i})^{2}-\operatorname{{\rm Var}}_{i}\bigg)\bigg\}=O(n^{2}/B^{3}),$
@@
$\mathbb{E}(M_{i}-A_{i})^{2}$ $=\frac{r^{2}}{n^{2}}\sum_{k=1}^{r-1}{{n-1}\choose{k}}\bigg(\frac{{{r-1}\choose{k}}}{{{n-1}\choose{k}}}-\frac{{{r}\choose{k+1}}}{{{n}\choose{k+1}}}\bigg)^{2}V_{k+1}+\frac{r^{2}}{n^{2}}\sum_{k=2}^{r}{{n-1}\choose{k}}\bigg(\frac{{{r-1}\choose{k}}}{{{n-1}\choose{k}}}-\frac{{{r}\choose{k}}}{{{n}\choose{k}}}\bigg)^{2}V_{k}$
$\precsim\frac{r^{2}(n-1)}{n^{2}}\bigg(\frac{{r-1}}{{n-1}}-\bigg(\frac{{\binom{r}{2}}}{{\binom{n}{2}}}\bigg)\bigg)^{2}\binom{r}{2}^{-1}\operatorname{{\rm Var}}[\widehat{f}^{b}(\mathbf{x}_{*})]$
$\precsim\frac{2r^{2}\operatorname{{\rm Var}}[\widehat{f}^{b}(\mathbf{x}_{*})]}{n^{3}}.$ (C.10)
Here, the first inequality is by the fact that summation in the first line is maximized in the second-order terms and $\sum_{k=1}^{r}{r\choose{k}}V_{k}=\operatorname{{\rm Var}}(\widehat{f}^{b}(\mathbf{x}_{*}))\leq C<+\infty$, and the second inequality is by simple calculation. We hence have
@@ Appendix D Proofs in Section C
$\mathbb{E}(Y|\eta)=\psi^{\prime}(\eta),\quad\operatorname{{\rm Var}}(Y|\eta)=\psi^{\prime\prime}(\eta)\geq 0.$
We conclude that $\psi(\cdot)$ is a convex function. By Assumption 3.1, the non-degeneracy and the existence of distinct values in the support of $p(y|\eta)=h(y)\exp(\eta y-\psi(\eta))$ ensure that $\operatorname{{\rm Var}}(Y|\eta)>0$ for all $\eta\in[-C,C]$. This guarantees that the second derivative of the log-partition function, $\psi^{\prime\prime}(\eta)=\operatorname{{\rm Var}}(Y|\eta)$, is strictly positive. Furthermore, since $\mathbb{E}(Y|\eta)$ and $\mathbb{E}(Y^{2}|\eta)$ are defined as smooth integrals with respect to $p(y|\eta)$, they are continuous in $\eta$. As a result, $\psi^{\prime\prime}(\eta)$ is also continuous. By the compactness of $\eta\in[-C,C]$, there exists a constant $c>0$ such that:
@@
Lemma D.1 establishes the differences between $R_{n}(\widehat{f}_{n},f_{0})$ and $\widehat{R}_{n}(\widehat{f}_{n},f_{0})$. We now propose the following lemma, which addresses the challenges arising from the dependency of $\varepsilon_{i}$ on the covariates $\mathbf{x}_{i}$. This approach is different from previous work, such as Schmidt-Hieber (2020) and Fan and Gu (2024), which assumed independence between $\varepsilon_{i}$ and $\mathbf{x}_{i}$.
@@
$\operatorname{{\rm Var}}\big(g_{j}(\bm{X}^{\prime}_{i},\mathbf{x}_{i})\big)$ $=2\operatorname{{\rm Var}}(\ell(\bm{X};f_{j},f_{0}))$
@@
$\mathbb{P}(|\sum_{i=1}^{n}\xi_{i}|\geq t_{0})\leq 2\exp\bigg(-\frac{t_{0}^{2}}{2Mt_{0}/3+2\sum_{i=1}^{n}\operatorname{{\rm Var}}(\xi_{i})}\bigg).$
@@
$\operatorname{{\rm Var}}(\xi_{j}|(\mathbf{x}_{i})_{i})$ $=\frac{\sum_{i=1}^{n}\operatorname{{\rm Var}}\big(\varepsilon_{i}(f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i}))|(\mathbf{x}_{i})_{i}\big)}{r_{j}^{2}}$
@@
where the first inequality is by $\operatorname{{\rm Var}}(U|X)\leq\mathbb{E}(U^{2}|X)$ for all random variables $X$ and $U$, and the second inequality is by the definition of $r_{j}=\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\vee\sqrt{\sum_{i=1}^{n}(f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i}))^{2}/n}$, and $\mathbb{E}\big(\varepsilon_{i}^{2}|(\mathbf{x}_{i})_{i}\big)\leq 2\kappa^{2}$. Also, letting $a_{ij}=[f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i})]/r_{j}$, we have